Here is the question:
The product of two numbers is -35, and of the numbers is 17 more than twice the other. Find all pairs of numbers that satisfy these conditions. The answer was given as -7/2, 10 : -5,7. Please explain how these answers were found!

Answer by Boreal(15235)

xy=-35, x and y being the numbers
y=2x+17
x(2x+17)=-35, use substitution
2x^2+17x+35=0
(2x+7)(x+5)=0
x=-7/2,-5.
Those are values for x, not the pair of numbers. With those values, you can get y.
When it is -7/2, y has to be +10
When x=-5, y has to be +7